In the city of Growville, the equilibrium employment is 100,000 workers and the equilibrium wage is $100 per day. The wage elasticity of demand for labor is −1.0 and the wage elasticity of supply of labor is 5.0. Suppose the demand for labor increases by 18 percent. Illustrate the effects of the increase in labor demand on the urban labor market, including values for the equilibrium wage and equilibrium total employment.
Answer #1

Increase in the demand for labor will shift the demand curve out

As a result, wage rate and employment both are increased

Increase in wages = % increase in demand / (wage elasticity of demand + wage elasticity of supply)

= 18% / (1 + 5)

= 3%

Hence, new wage rate is 100 + 3%*100 = $103

New employment = old employment - wage elasticity of demand*% increase in wages*old employment + wage elasticity of supply*% increase in wages*old employment

= 100000 - 1*3%*100000 + 5*3%*100000 = 112000

This new equilibrium is shown at F.